lim {(integration of F(t) dt) 4m 2 to sec^2 x} divided by {x^2- pi^2/16}


 x->pi/4

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans: 0
Sol:
L = \lim_{x\rightarrow \frac{\pi }{4}}\frac{\int_{2}^{sec^{2}x}F(x)dx}{x^{2}-\frac{\pi ^{2}}{16}}
It is zero by zero form. So apply L’hospital rule.
L = \lim_{x\rightarrow \frac{\pi }{4}}\frac{F(sec^{2}x)-F(2)}{2x} = \frac{F(2)-F(2)}{2.\frac{\pi }{4}}
L = 0
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty